OpenStudy (anonymous):

divide and simplify b+2/b-3divided by 2b+4/b-2

OpenStudy (saifoo.khan):

b-2 divide by 2(b-3)

OpenStudy (anonymous):

\[\frac{b+2}{b-3}\div\frac{2b+4}{b-2}\]right?

OpenStudy (saifoo.khan):

yes

OpenStudy (anonymous):

time and flip \[\frac{b+2}{b-3}\times\frac{b-2}{2b+4}\]

OpenStudy (anonymous):

fator you get \[\frac{b+2}{b-3}\times\frac{b-2}{2(b+2)}\]b+2 cancle\[\frac{b-2}{2(b-3)}\]